Use integration by parts to integrate x sec x tan x dx ∫
(2)
b) Integrate 3( x )1 4x 12x 5 dx 2

+ + + (3)
c) Integrate dx
x( x()4 )1
x x 5
1
Expert's answer
2020-06-30T18:06:18-0400

"(1)
\int xsecx tanx dx"

Using integration by parts we get

"\smallint udv=uv-\smallint vdu"

u=x dv= secxtanxdx

du= dx v= secx


"\smallint xsecxtanxdx= xsecx-\smallint secxdx 

\implies xsecx- In |secx+ tanx|+C"

(2)

"I=\int (3x+1)\sqrt{(2x+3)^2-4}dx=\frac{1}{2}\int(\frac{3(u-3)}{2}+1)\sqrt{u^2-4}du\,

where \,u=2x+3 \,hence \,I=\frac{3}{4}\int u\sqrt{u^2-4} du-\frac{7}{4}\int\sqrt{u^2-4} du\,\,
if\, u=\frac{2}{ \ cost},du=\frac{2\ sint}{ \cos^2t}dt\,\,then \,I=3\int\frac{ \sin^3t}{ \cos^4t} dt-\frac{7}{2}\int \frac{ \sin^2t}{\cos^3t} dt\,
as H.B.Dwight \,Tables\, of \,integrals , \,N.Y.,\,The \,Mcmillan\, corp,\,1961,\,f.(452.23),\,f.(452.39) \,hence\,\,

I=\frac{1-3\cos^2t}{ \cos^3t}-\frac{7}{2}(\frac{ \ sint}{2\ cos^2t}-\frac{1}{2}\ln\tan(\frac{\pi}{4}+\frac{t}{2}))\,\,where \,t=\arccos\frac{2}{2x+3}"

(3)


"I=\int(\frac{1}{x+1}+\frac{1}{x^2+4}) dx= \int ln |x+1 |+\frac{1}{2}\arctan\frac{x}{2}+C"